What Z-scores correspond to separating middle 70% of the normal distribution from the rest of the distribution?
Answer #1

Solution :

Solution :
P(-z < Z < z) = 0.70

P(Z < z) - P(Z < z) = 0.70

2P(Z < z) - 1 = 0.70

2P(Z < z) = 1 + 0.70

2P(Z < z) = 1.70

P(Z < z) = 1.70 / 2

P(Z < 1.04) = 0.85

z = 1.04

P(-1.04 < Z < 1.04) = 0.70

Z-scores : -1.04 , +1.04